OpenStudy (anonymous):

show that all rectangles with perimeter P the square has the greatest area. i dont understand...

OpenStudy (anonymous):

is this algebra or calculus?

OpenStudy (anonymous):

its algebra, but it says precalculus honors on the worksheet. im in the quadratics unit

OpenStudy (anonymous):

yes, you can solve it using quadratics

OpenStudy (anonymous):

lets say the perimeter is 100 and call one side x and the other y 

OpenStudy (anonymous):

then \[2x+2=y=100\\ A=xy\] if you solve the first equation for \(y\) you get \(y=50-x\) so \[A(x)=x(50-x)=50x-x^2\]

OpenStudy (anonymous):

damn typo first line should be \[2x+2y=100\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the quadratic is therefore \[A(x)=50x-x^2\] and the max is at the vertex first coordinate of the vertex is always \(-\frac{b}{2a}\) which in this case is \(-\frac{50}{2\times(-1)}=25\)

OpenStudy (anonymous):

so the if the perimeter is 100, then the max is when \(x=25\) i.e. each side is 25 and it is a square

OpenStudy (anonymous):

to prove it for all perimeters, replace the \(100\) i used by a \(P\) and repeat the process

OpenStudy (anonymous):

so it's simply proving that for any given perimeter of a rectangle, the square's area will be greater than the rectangle?